What power of spectacle lens is needed to correct the vision of a nearsighted person whose far point is 30.0cm. Assume the spectacle lens is held 1.50cm away from the eye by eyeglass frames.


1
Expert's answer
2021-03-18T14:01:48-0400

Given,

Far point of the person = 30cm

Distance between eye and eyeglass frame =1.5 cm

So, image distance "(v) = -(30cm-1.5cm) =- 28.5cm"

Here the distance is negative because it is in same side of the spectacle as the object.

Object distance "(u)=\infty"

We know that,

"\frac{1}{f}=\frac{1}{v}-\frac{1}{u}"

Now, substituting the values,

"\Rightarrow \frac{1}{f}=\frac{1}{-28.5}-\frac{1}{\infty}"


"\Rightarrow \frac{1}{f}=\frac{1}{-28.5}"


We know that,

Power of lens "(P)=\frac{100}{f}"


"\Rightarrow P=\frac{100}{-28.5}D"


"\Rightarrow P =-3.5D"

Here, negative sign representing that the lens should be concave.